A Bibliography of Publications of M. Victor Wickerhauser

M. Victor Wickerhauser Washington University

Department of Mathematics Campus Box 1146, One Brookings Drive

Saint Louis, Missouri 63130 USA

Tel: +1 314 935 6771, +1 314 935 6771 (secretary) FAX: +1 314 935 5799

E-mail: Victor@Math.WUStL.Edu (Internet)

17 November 2023 Version 1.12

Abstract

This bibliography records publications of M. Victor Wickerhauser.

Title word cross-reference

2xx [Wic81].

= [oM90].

Acoustic [Wic92a]. Adapted [CW90a, CMW92b, CW92b, CW93, CW94, WW92, WW93, Wic92b, Wic93b, Wic94b, Wic94a, WFG^{+}94, Wic93a]. Algorithm [Wic81]. Algorithms [CMP94, CW92a, Wic91b, Wic94h, Wic94f]. Analysis [CMW91, CMW92b, CMW92d, CW92b, CW93, CW94, FFW95, HNW95, MR93, Wic91a, Wic92b, Wic94b, Wic94d]. Application [GWF94a, GWF94b].

Applications [BF92, BBHB94, Chu92, CMP94, CMW92b, LU94, Mey92, MR93, MJ94, RBC^{+}92, Szu94, Wic94h]. Applied [O’M92]. Appliqués [Lyo91]. Approximate [Wic91a, Wic94d, Wic94f]. Approximations [Wic94g, Wic93d]. Atoms [Wic93b]. Automatic [MM94].

Band [Wic90b]. Based [CW92a]. Bases [AWW92, CW90a, Wic93a, Wic93c, WFG^{+}94]. Basis [CW92a, Wic90b]. Benjamin [CW90b]. Best [CW90a, CW92a, Wic90b, Wic93a]. Best-Adapted [CW90a, Wic93a]. Best-Basis [Wic90b]. Burgers [HPW92].

Calculus [Wic94a]. Carolina [CPBE94]. Centenary [CPBE94]. Chaotic [Tat84]. Chemical [MJ94]. Classification [Wic94f]. Coding [Wic90b, Wic94c]. Coefficients [PW94]. Coifman [AWW92]. Comparison [Wic94c, WFG^{+}94]. Complicated [Wic94d, Wic94g, Wic93d]. Component [Wic94d]. Compression [CMQW90, CMQW93, CMQW94, FGM^{+}92, Wic90b, Wic92c, Wic94c, WFG^{+}94, ZSW91, Wic92a]. Computation [Wic93b]. Computer [ST91]. computers [RW92]. Conference [O’M92]. Connection [PW94]. Construction [AWW92]. Cornelius [CPBE94]. Cosine [AWW92, Wic94c, WFG^{+}94]. Curves [ST91]. Cyber [Wic81].

Data [ZSW91]. December [CPBE94]. Decomposition [HPW92]. Denoise [CMW92a]. Denoising [CW94]. Department [pasnt]. Different [Dau93]. Dimensional [FGM^{+}92, GWF94a, GWF94b, WFG^{+}94]. Discrimination [Wic94d]. D’Ondes [Lyo91].

Efficiency [WFG^{+}94]. Elementary [GW88]. Elias [FFW95]. Engineering [MJ94]. Entropy [CW92a]. Equation [CW90b, Wic88, HPW92]. Essays [FFW95]. Evolutions [Wic85, Wic87]. Extraction [CW94].

Factor [Wic91a]. Factored [Wic81]. Fast [Wic91a, Wic94f]. Feature [CW94, Wic94d]. Flow [WFG^{+}94]. Flows [FGM^{+}92]. Form [Wic88]. Fourier [FFW95]. Frequency [CGT89, HNW95, Wic93b]. ftp.c3 [Lab93]. Functional [Wic94a].

Geophysics [FGK94]. Graphics [ST91].

Hamilton [Wic88]. Harmonic [CMW91]. Heat [Wic85, Wic87]. High [Wic92c]. High-Resolution [Wic92c]. Honor [FFW95]. Humans [MM94].

ICIAM [O’M92]. Identification [MM94]. II [HJ93, LU94, ST91]. Image [LU94, Wic94f]. Images [ZSW91]. Implicit [Wic81]. Improved [FGM^{+}92]. Industrial [O’M92]. Information [HJ93]. INRIA [Wic91b]. Inspection [MM94]. International [O’M92]. Inverse [Wic87]. Inversion [Wic94d, Wic94g, Wic93d].

Jacobians [Wic93d, Wic94g]. July [O’M92].

Kadomtsev [Wic88].

Lanczos [CPBE94]. Large [Wic94d]. Large-rank [Wic94d]. Lectures [CW91, Wic91b]. Library [Wic92b]. Linéaires [Lyo91]. Local [AWW92, WW92, WW93, Wic94c, WFG^{+}94]. Localization [Wic94e]. Localized [Wic93c]. Louis [wua91].

M [FFW95]. Malvar [DFWW92, WW94]. Maps [Wic94d, Wic94g, Wic93d]. Marietta [CW91]. Martin [CW91]. Mathematics [BF92, O’M92, pasnt]. Matrix [GWF94a, GWF94b, Wic90a]. Matrix-Vector [GWF94a, GWF94b]. Methods [CGT89, Wic94c]. Meyer [AWW92, Wic94h]. Modeling [CW94]. Multiplication [GWF94a, GWF94b, PW94, Wic90a].

Navier [Wic81]. NeXT [RW92]. Non [Lyo91]. Non-Linéaires [Lyo91]. Nonlinear [Wic85]. Nonstandard [Wic90a]. North [CPBE94]. Numerical [CMW91, CW93].

Ondelettes [Lyo91, WW94]. Ono [CW90b]. Operator [Wic85, Wic87]. Orthonormal [Wic93c].

Packet

[CW90a, FGM^{+}92, GWF94a, GWF94b, Wic91b, Wic93a, Wic94c, WFG^{+}94]. Packets [CMQW90, CMW92b, CMW92c, CMQW93, CMQW94, ZSW91, Wic92a].

Paquets [Lyo91]. Parallel [GWF94a, GWF94b]. Parole [WW94]. Performance [Wic81]. Perspectives [Dau93]. Petviashvili [Wic88]. Phase [CGT89]. Phenomena [Tat84]. Picture [Wic90b, Wic92c, Wic94c]. PIS [Tat84]. Predictability [FGM^{+}92]. Principal [Wic94d]. Problèmes [Lyo91]. Proceedings [CPBE94, O’M92, Tat84]. Processing [CMQW90, CMW92d, CMQW93, CW93, CMQW94, HJ93, LU94, WW92, WW93, Wic94f]. Progress [MR93]. Properties [CMW92c].

Raleigh [CPBE94]. rank [Wic94d]. Recognition [Wic94f]. Resolution [Wic92c].

Scale [HPW92]. Scattering [CW90b, Wic87]. Second [O’M92]. Segmentation [DFWW92]. Selection [CW92a]. Series [PW94]. Short [PW94]. Signal [DFWW92, CMQW90, CMW92d, CMQW93, CW93, CMQW94, LU94, Wic94d, Wic92a]. Sine [AWW92]. Size [CMW92c]. Smooth [AWW92, Wic93c]. Software [Wic94b]. Space [CGT89]. Speech [DFWW92, WW92, WW93]. St [wua91]. Stein [FFW95]. Still [Wic92c]. Stokes [Wic81]. Sub [Wic90b]. Sub-Band [Wic90b]. Surfaces [ST91]. Symposium [Tat84]. Systems [MM94].

Tatsumi [Tat84]. Techniques [Wic94e]. Their [BBHB94, RBC^{+}92]. Theory [Chu92, CMP94, Wic94b]. Time [CGT89, HNW95, Wic93b, Wic94e]. Time-Frequency [CGT89, HNW95, Wic93b]. Tool [CW94]. Toolkit [CW93]. Traitement [WW94]. Transform [CW90b, GWF94a, GWF94b, WW92, WW93, Wic94c]. Transforms [Wic94e]. Trigonometric [WW92, WW93].

Turbulence [Tat84, ZSW91]. Turbulent [FGM^{+}92, WFG^{+}94]. Tutorial [Chu92]. Two [FGM^{+}92, GWF94a, GWF94b, WFG^{+}94, Wic94f]. Two-Dimensional [FGM^{+}92, WFG^{+}94].

University [pasnt, wua91]. Using

[FGM^{+}92, PW94, ZSW91].

v2.0 [Wic92b]. Variables [Wic87]. Vector [GWF94a, GWF94b]. version [RW92]. via [DFWW92]. Vision [ST91]. Visual [HJ93].

Washington [wua91]. Wave [CW90a]. Waveform [CMW92b, CW92b, CW93, CW94, Wic92b, Wic94a]. Wavelet [CMQW90, CW91, CMW92b, CMW92c, CMW92d, CMQW93, CMQW94, FGM^{+}92, GWF94a, GWF94b, LU94, MR93, PW94, Szu94, Wic91b, Wic93a, Wic93d, Wic94b, Wic94c, WFG^{+}94, Wic94f, Wic94e, Wic94g, ZSW91, Wic92a]. Wavelet-Packets [CMW92b]. Wavelets [AWW92, BF92, BBHB94, DFWW92, Chu92, CMP94, CW92b, CW93, CGT89, Dau93, FGK94, GW88, HNW95, Mey92, MJ94, RBC^{+}92, Wic94h, Wic94d]. WPLab [RW92].

Yale [pasnt]. Yves [Wic94h].

References

Auscher:1992:LSC

[AWW92] Pascal Auscher, Guido Weiss, and Mladen Victor Wickerhauser. Local sine and cosine bases of Coifman and Meyer and the construction of smooth wavelets. In Chui [Chu92], pages 237–256. ISBN 0-12-174590-2. URL ftp:

                                                   //wuarchive.wustl.edu/doc/

techreports/wustl.edu/math/
papers/aww.ps.Z.

Byrnes:1994:WTA

[BBHB94] James S. Byrnes, Jennifer L.

Byrnes, Kathryn A. Hargreaves, and Karl Berry, editors. Wavelets and Their Applications, volume 442 of NATO ASI Series C: Mathematical and Physical Sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-3078-1. Proceedings of the NATO Advanced Study Institute at Il Ciocco, Barga, Italy in August, 1992.

Benedetto:1992:WMA

[BF92] John J. Benedetto and Michael Frazier, editors. Wavelets: Mathematics and Applications. Studies in Advanced Mathematics. CRC Press, 2000 Corporate Blvd., Boca Raton, FL 33431, USA, 1992. ISBN 0-8493-8271-8.

Combes:1989:WTF

[CGT89] Jean-Michel Combes, Alexander Grossmann, and Philippe Tchamitchian, editors. Wavelets: Time-Frequency Methods and Phase Space. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, second edition, 1989. ISBN 0-387-53014-2.

Chui:1992:WTT

[Chu92] Charles K. Chui, editor. Wavelets– A Tutorial in Theory and Applications. Academic Press, New York, NY, USA, 1992. ISBN 0-12-174590-2.

Chui:1994:WTA

[CMP94] Charles K. Chui, Laura Montefusco, and Luigia Puccio, editors. Wavelets: Theory, Algorithms, and Applications, Proceedings of the International Conference in Chui:1994:WTA, Sicily, 14– 20 October 1993. University of Messina, Academic Press, New York, NY, USA, 1994. ISBN 0-12-174575-9.

Coifman:1990:SPC

[CMQW90] Ronald R. Coifman, Yves Meyer, Stephen R. Quake, and Mladen Victor Wickerhauser. Signal processing and compression with wavelet packets. Preprint, Yale University, New Haven, April 1990. URL ftp://

math.yale.edu/pub/wavelets/
cmqw.tex.

Coifman:1993:SPC

[CMQW93] Ronald R. Coifman, Yves Meyer, Stephen R. Quake, and Mladen Victor Wickerhauser. Signal processing and compression with wavelet packets. In Meyer and Roques [MR93], pages 77–93. ISBN 2-86332-130-

7. URL ftp://math.yale.edu/
pub/wavelets/cmqw.tex.

Coifman:1994:SPC

[CMQW94] Ronald R. Coifman, Yves Meyer, Stephen R. Quake, and Mladen Victor Wickerhauser. Signal processing and compression with wavelet packets. In Byrnes et al. [BBHB94], pages 363–379. ISBN 0-7923-3078-

1. URL ftp://math.yale.edu/

pub/wavelets/cmqw.tex. Pro-

ceedings of the NATO Advanced Study Institute at Il Ciocco, Barga, Italy in August, 1992.

Coifman:1991:NHA

[CMW91] Ronald R. Coifman, Yves Meyer, and Mladen Victor Wickerhauser. Numerical harmonic analysis. In Fefferman et al. [FFW95], pages 162–174. ISBN 0-691-08655-

9.  URL ftp://math.yale.
edu/pub/wavelets/numharan.

tex. Proceedings of the Princeton Conference in Harmonic Analysis, held 13–17 May 1991.

Coifman:1992:D

[CMW92a] Ronald R. Coifman, Fazal Majid, and Mladen Victor Wickerhauser. Denoise. Available from Fast Mathematical Algorithms and Hardware Corporation, 1020 Sherman Ave., Hamden, CT 06514 USA, 1992.

Coifman:1992:AWA

[CMW92b] Ronald R. Coifman, Yves Meyer, and Mladen Victor Wickerhauser. Adapted waveform analysis, wavelet-packets and applications. In O’Malley, Jr. [O’M92], pages 41–50. ISBN 0-89871-302-

1. URL ftp://math.yale.edu/
pub/wavelets/adaptwave1.tex.

Coifman:1992:SPW

[CMW92c] Ronald R. Coifman, Yves Meyer, and Mladen Victor Wickerhauser. Size properties of wavelet packets. In Ruskai et al. [RBC^{+}92], pages 453–470. ISBN 0-86720-

225-4. URL ftp://math.yale.

                                        edu/pub/wavelets/sizeprop.
                                        tex.

Coifman:1992:WAS

[CMW92d] Ronald R. Coifman, Yves Meyer, and Mladen Victor Wickerhauser. Wavelet analysis and signal processing. In Ruskai et al. [RBC^{+}92], pages 153–178. ISBN 0-86720-225-4. URL ftp:

                                        //wuarchive.wustl.edu/doc/
                                        techreports/wustl.edu/math/
                                        papers/wasp.ps.Z.

Chu:1994:PCL

[CPBE94] Moody Chu, Robert Plemmons, David Brown, and Donald Ellison, editors. Proceedings of the Cornelius Lanczos Centenary, Raleigh, North Carolina, 12–17 December 1993. SIAM, SIAM, Philadelphia, PA, USA, 1994. ISBN 0-89871-339-0.

Coifman:1990:BAW

[CW90a] Ronald R. Coifman and Mladen Victor Wickerhauser. Best-adapted wave packet bases. Preprint, Yale University, New Haven,

                                        1990. URL ftp://math.yale.
                                        edu/pub/wavelets/bestbase.
                                        tex.

Coifman:1990:STB

[CW90b] Ronald R. Coifman and Mladen Victor Wickerhauser. The scattering transform for the Benjamin– Ono equation. Inverse Problems, 6:825–861, 1990. CO-DEN INPEEY. ISSN 0266-5611 (print), 1361-6420 (elec-

                                        tronic). URL ftp://math.yale.
                                        edu/pub/papers/benj-ono.tex.

Coifman:1991:MMW

[CW91] Ronald R. Coifman and Mladen Victor Wickerhauser. Martin– Marietta Wavelet Lectures. Martin– Marietta Corporation, 21–25 October 1991.

Coifman:1992:EBA

[CW92a] Ronald R. Coifman and Mladen Victor Wickerhauser. Entropy based algorithms for best basis selection. IEEE Transactions on Information Theory, 32:712–718, March 1992. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic). URL ftp://

math.yale.edu/pub/wavelets/
entropy.tex.

Coifman:1992:WAW

[CW92b] Ronald R. Coifman and Mladen Victor Wickerhauser. Wavelets and adapted waveform analysis. In Benedetto and Frazier [BF92], pages 399–423. ISBN 0-8493-

8271-8. URL ftp://wuarchive.
wustl.edu/doc/techreports/
wustl.edu/math/papers/wawa.
ps.Z.

Coifman:1993:WAW

[CW93] Ronald R. Coifman and Mladen Victor Wickerhauser. Wavelets and adapted waveform analysis: A toolkit for signal processing and numerical analysis. In Daubechies [Dau93], pages 119–153. ISBN 0-8218-

5503-4. URL ftp://wuarchive.
wustl.edu/doc/techreports/
wustl.edu/math/papers/wawa.

ps.Z. Minicourse lecture notes.

Coifman:1994:AWA

[CW94] Ronald R. Coifman and Mladen Victor Wickerhauser. Adapted waveform analysis as a tool for modeling, feature extraction, and denoising. Optical Engineering, 33 (7):2170–2174, July 1994. CO-DEN OPEGAR. ISSN 0091-3286 (print), 1560-2303 (electronic).

                                        URL ftp://wuarchive.wustl.
                                        edu/doc/techreports/wustl.
                                        edu/math/papers/awaatr.ps.Z.

Special issue on Adapted Wavelet Analysis.

Daubechies:1993:DPW

[Dau93] Ingrid Daubechies, editor. Different Perspectives on Wavelets, number 47 in Proceedings of Symposia in Applied Mathematics. Amer. Math. Soc., Providence, RI, USA, 11-12 January 1993. ISBN 0-8218-5503-4.

ChristopheAlessandro:1992:SSS

[DFWW92] Christophe D’Alessandro, Xiang Fang, Eva Wesfreid, and Mladen Victor Wickerhauser. Speech signal segmentation via Malvar wavelets. In Meyer and Roques [MR93], pages 305–308. ISBN 2-86332-130-7.

Fefferman:1995:EFA

[FFW95] Charles Fefferman, Robert Fefferman, and Stephen Wainger, editors. Essays on Fourier Analysis in Honor of Elias M. Stein. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-08655-9. Proceedings of the Princeton Conference in Harmonic Analysis, held 13–17 May 1991.

Foufoula-Georgiou:1994:WG

[FGK94] Efi Foufoula-Georgiou and Praveen Kumar, editors. Wavelets in Geophysics. Academic Press, New York, NY, USA, 1994. ISBN 0-12-262850-0.

Farge:1992:IPT

[FGM^{+}92] Marie Farge, Eric Goirand, Yves Meyer, Frédéric Pascal, and Mladen Victor Wickerhauser. Improved predictability of twodimensional turbulent flows using wavelet packet compression. Fluid Dynamics Research, 10: 229–250, 1992. URL ftp:

//wuarchive.wustl.edu/doc/
techreports/wustl.edu/math/
papers/fdr92.ps.Z.

Gootman:1987:EW

[GW88] Elliot Gootman and Mladen Victor Wickerhauser. Elementary wavelets. Preprint, 02021-88, Mathematical Sciences Research Institute, Berkeley, California, 1987-88. URL ftp://

math.yale.edu/pub/wavelets/
elemwave.tex.

Goirand:1994:PTDa

[GWF94a] Eric Goirand, Mladen Victor Wickerhauser, and Marie Farge. A parallel two dimensional wavelet packet transform and its application to matrixvector multiplication. Preprint, Washington University Department of Mathematics, St. Louis, Missouri, 1994. URL ftp:

                                        //wuarchive.wustl.edu/doc/
                                        techreports/wustl.edu/math/
                                        papers/parallel.zip.

Goirand:1994:PTDb

[GWF94b] Eric Goirand, Mladen Victor Wickerhauser, and Marie Farge. A parallel two dimensional wavelet packet transform and its application to matrix-vector multiplication. In Motard and Joseph [MJ94], pages 275–319. ISBN 0-7923-9461-5. URL ftp:

                                        //wuarchive.wustl.edu/doc/
                                        techreports/wustl.edu/math/
                                        papers/parallel.zip.

Huck:1993:VIP

[HJ93] Friedrich O. Huck and Richard D. Juday, editors. Visual Information Processing II, volume 1961 of SPIE Proceedings. SPIE, Orlando, Florida, April 1993. ISBN 0-8194-1197-3.

Hess-Nielsen:1995:WTF

[HNW95] Nikolaj Hess-Nielsen and Mladen Victor Wickerhauser. Wavelets and time-frequency analysis. Proceedings of the IEEE, 1995. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (elec-

                                        tronic). URL ftp://wuarchive.
                                        wustl.edu/doc/techreports/
                                        wustl.edu/math/papers/wtfa.

ps.Z. To appear in the special issue on wavelet applications.

Heurtaux:1992:SDB

[HPW92] Frédéric Heurtaux, Fabrice Planchon, and Mladen Victor Wickerhauser. Scale decomposition in Burgers’ equation. In Benedetto and Frazier [BF92], pages 505– 523. ISBN 0-8493-8271-8. URL

ftp://wuarchive.wustl.edu/
doc/techreports/wustl.edu/
math/papers/burgers.ps.Z.

Laboratory:1993:FC

[Lab93] Los Alamos National Laboratory.

ftp.c3. InterNet Anonymous File Transfer (ftp) Site ftp.c3.lanl.

gov [128.165.21.64], 1993.

Laine:1994:WAS

[LU94] Andrew Laine and Michael A.

Unser, editors. Wavelet Applications in Signal and Image Processing II, volume 2303. SPIE, SPIE, 27–29 July 1994. ISBN 0-8194-1627-4.

Lyons:1991:PNL

[Lyo91] Pierre-Louis Lyons, editor. Problèmes Non-Linéaires Appliqués, Ondelettes et Paquets D’Ondes. IN-RIA, Roquencourt, France, 17– 21June 1991. Minicourse lecture notes.

Meyer:1992:WA

[Mey92] Yves Meyer, editor. Wavelets and Applications, Proceedings of the International Conference “Wavelets and Applications” Marseille, May 1989. LMA/CNRS, Masson, Paris, France, 1992.

Motard:1994:WAC

[MJ94] Rodolphe L. Motard and Babu Joseph, editors. Wavelets Applications in Chemical Engineering. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-9461-5.

Mammone:1994:ASI

[MM94] Richard J. Mammone and J. David Murley, Jr., editors. Automatic Systems for the Identification and Inspection of Humans, volume 2277 of SPIE Proceedings. SPIE, San Diego, California, 24– 29 July 1994.

Meyer:1993:PWA

[MR93] Yves Meyer and Sylvie Roques, editors. Progress in Wavelet Analysis and Applications, Proceedings of the International Conference “Wavelets and Applications,” Toulouse, France, 8– 13 June 1992. Observatoire Midi-Pyrénées de l’Université Paul Sabatier, Editions Frontières, Gif-sur-Yvette, France, 1993. ISBN 2-86332-130-7.

Mathematics:1990:PYY

[oM90] Yale University Department of Mathematics. yalevm.ycc.

yale.edu [192.31.2.2]. In-

stant Math Preprints Abstracts Database, 1990.

OMalley:1992:IPS

[O’M92] Robert E. O’Malley, Jr., editor. ICIAM 91: Proceedings of the Second International Conference on Industrial and Applied Mathematics, 8–12 July, 1991. SIAM, SIAM, Philadelphia, PA, USA, 1992. ISBN 0-89871-302-1.

pascal:1989:YUD

[pasnt] pascal. Yale University Department of Mathematics. InterNet Anonymous File Transfer (ftp)

Site pascal.math.yale.edu  [

128.36.23.1], 1989–present.

Perrier:1994:MSW

[PW94] Valerie Perrier and Mladen Victor Wickerhauser. Multiplication of short wavelet series using connection coefficients. Technical report, Washington University, Saint Louis, Missouri 63130, 1994. 15 pp.

Ruskai:1992:WTA

[RBC^{+}92] Mary Beth Ruskai, Gregory Beylkin, Ronald Coifman, Ingrid Daubechies, Stéphane Mallat, Yves Meyer, and Louise Raphael, editors. Wavelets and Their Applications. Jones and Bartlett, Boston, 1992. ISBN 0-86720-225-4.

Rochberg:1992:WVN

[RW92] David Rochberg and Mladen Victor Wickerhauser. WPLab version 3.03 (for NeXT computers). Available by anonymous ftp from [wua91], 1992. URL ftp:

//wuarchive.wustl.edu/doc/
techreports/wustl.edu/math/
software/WPLab3.03.tar.Z.

Silbermann:1991:CSC

[ST91] Martine J. Silbermann and Hemant D. Tagare, editors. Curves and Surfaces in Computer Vision and Graphics II, volume 1610 of SPIE Proceedings. SPIE, Boston, October 1991. ISBN 0-8194-0747-X.

Szu:1994:WA

[Szu94] Harold H. Szu, editor. Wavelet Applications, volume 2242 of SPIE Proceedings. SPIE, Orlando, Florida, 5–8 April 1994. ISBN 0-8194-1546-4.

Tatsumi:1984:PIS

[Tat84] T. Tatsumi, editor. Proceedings of the Tatsumi:1984:PIS Symposium on Turbulence and Chaotic Phenomena. North-Holland, Amsterdam, The Netherlands, 1984.

Wickerhauser:1994:ECW

[WFG^{+}94] Mladen Victor Wickerhauser, Marie Farge, Eric Goirand, Eva Wesfreid, and Echeyde Cubillo. Efficiency comparison of wavelet packet and adapted local cosine bases for compression of a two-dimensional turbulent flow. In Chui et al. [CMP94]. ISBN 0-12-174575-9. URL ftp:

                                        //wuarchive.wustl.edu/doc/
                                        techreports/wustl.edu/math/
                                        papers/taormina2.ps.Z.

Wickerhauser:1981:CPI

[Wic81] Mladen Victor Wickerhauser. Cyber 2xx performance on an implicit factored Navier–Stokes algorithm. Preprint, NASA Ames Research Center, 1981.

Wickerhauser:1985:NEH

[Wic85] Mladen Victor Wickerhauser. Nonlinear Evolutions of the Heat Operator. PhD thesis, Yale University, New Haven, Connecticut, May 1985.

Wickerhauser:1987:ISH

[Wic87] Mladen Victor Wickerhauser. Inverse scattering for the heat operator and evolutions in 2+1 variables. Communications in Mathematical Physics, 108:67–89, 1987. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic).

Wickerhauser:1988:HFK

[Wic88] Mladen Victor Wickerhauser.

Hamilton’s form for the Kadomtsev– Petviashvili equation. Journal of Mathematical Physics, 29:2300– 2302, 1988. CODEN JMAPAQ. ISSN 0022-2488.

Wickerhauser:1990:NMM

[Wic90a] Mladen Victor Wickerhauser.

Nonstandard matrix multiplication. Preprint, Yale University, 15 May 1990. URL ftp://

math.yale.edu/pub/wavelets/
nsmatrix.tex.

Wickerhauser:1990:PCB

[Wic90b] Mladen Victor Wickerhauser.

Picture compression by best-basis sub-band coding. Preprint, Yale University, 1990. URL ftp://

math.yale.edu/pub/wavelets/
pic.tar.

Wickerhauser:1991:FAF

[Wic91a] Mladen Victor Wickerhauser.

Fast approximate factor analysis. In Silbermann and Tagare [ST91], pages 23–32. ISBN 0-8194-0747-X. URL ftp:

//wuarchive.wustl.edu/doc/
techreports/wustl.edu/math/
papers/fakle.ps.Z.

Wickerhauser:1991:ILW

[Wic91b] Mladen Victor Wickerhauser. IN-RIA lectures on wavelet packet algorithms. In Lyons [Lyo91], pages 31–99. URL ftp:

                                        //wuarchive.wustl.edu/doc/
                                        techreports/wustl.edu/math/
                                        papers/inria300.ps.Z. Mini-

course lecture notes.

Wickerhauser:1992:ASC

[Wic92a] Mladen Victor Wickerhauser. Acoustic signal compression with wavelet packets. In Chui [Chu92], pages 679–700. ISBN 0-12-174590-2. URL ftp://

                                        math.yale.edu/pub/wavelets/
                                        acoustic.tex.

Wickerhauser:1992:AWA

[Wic92b] Mladen Victor Wickerhauser. Adapted Waveform Analysis Library, v2.0. Fast Mathematical Algorithms and Hardware Corporation, Hamden, Connecticut, June 1992. Software Documentation.

Wickerhauser:1992:HRS

[Wic92c] Mladen Victor Wickerhauser. High-resolution still picture compression. Digital Signal Processing: a Review Journal, 2(4):204– 226, October 1992. URL ftp:

                                        //wuarchive.wustl.edu/doc/
                                        techreports/wustl.edu/math/
                                        papers/dsp.ps.Z.

Wickerhauser:1993:BAW

[Wic93a] Mladen Victor Wickerhauser. Best-adapted wavelet packet bases. In Daubechies [Dau93], pages 155–171. ISBN 0-8218-5503-4. Minicourse lecture notes.

Wickerhauser:1993:CAT

[Wic93b] Mladen Victor Wickerhauser.

Computation with adapted timefrequency atoms. In Meyer and Roques [MR93], pages 175–184. ISBN 2-86332-130-7. URL ftp:

//wuarchive.wustl.edu/doc/
techreports/wustl.edu/math/
papers/cwatfa.ps.Z.

Wickerhauser:1993:SLO

[Wic93c] Mladen Victor Wickerhauser.

Smooth localized orthonormal bases. Comptes rendus de l’Académie des sciences, Paris, 316:423–427, 1993. URL ftp:

//wuarchive.wustl.edu/doc/
techreports/wustl.edu/math/
papers/crasslob.ps.Z.

Wickerhauser:1993:WAJ

[Wic93d] Mladen Victor Wickerhauser.

Wavelet approximations to Jacobians and the inversion of complicated maps. In Proceedings of Math-Chem-Comp 1993. Rudjer Bošković Institute, Zagreb, Rovinj, Croatia, 21–25 June 1993.

URL ftp://wuarchive.wustl.
edu/doc/techreports/wustl.
edu/math/papers/rovinj.ps.Z.

Wickerhauser:1994:AWF

[Wic94a] Mladen Victor Wickerhauser. An adapted waveform functional calculus. In Chu et al. [CPBE94], pages 418–421. ISBN 0-89871-339-0.

Wickerhauser:1994:AWA

[Wic94b] Mladen Victor Wickerhauser. Adapted Wavelet Analysis from Theory to Software. A. K. Peters, Ltd., Wellesley, MA, USA, 9 May 1994. ISBN 1-56881-041-5. 486+xii pages, with diskette.

Wickerhauser:1994:CPC

[Wic94c] Mladen Victor Wickerhauser. Comparison of picture compression methods: Wavelet, wavelet packet, and local cosine transform coding. In Chui et al. [CMP94]. ISBN 0-12-174575-9.

Wickerhauser:1994:LRA

[Wic94d] Mladen Victor Wickerhauser. Large-rank approximate principal component analysis with wavelets for signal feature discrimination and the inversion of complicated maps. Journal of Chemical Information and Computer Science, 34(5):1036– 1046, September/October 1994.

URL ftp://wuarchive.wustl.
edu/doc/techreports/wustl.
edu/math/papers/rovinj.ps.

Z. Proceedings of Math-Chem-Comp 1993, Rovinj, Croatia.

Wickerhauser:1994:TLT

[Wic94e] Mladen Victor Wickerhauser. Time localization techniques for wavelet transforms. In Mammone and Murley, Jr. [MM94], pages vii + 290. URL ftp:

//wuarchive.wustl.edu/doc/
techreports/wustl.edu/math/
papers/timeloc.ps.Z.

Wickerhauser:1994:TFA

[Wic94f] Mladen Victor Wickerhauser.

Two fast approximate wavelet algorithms for image processing, classification, and recognition. Optical Engineering, 33(7):2225– 2235, July 1994. CODEN OPE-GAR. ISSN 0091-3286 (print), 1560-2303 (electronic). URL

ftp://wuarchive.wustl.edu/
doc/techreports/wustl.edu/
math/papers/oe.ps.Z. Special

issue on Adapted Wavelet Analysis.

Wickerhauser:1994:WAJ

[Wic94g] Mladen Victor Wickerhauser.

Wavelet approximations to Jacobians and the inversion of complicated maps. In Szu [Szu94], pages 100–118. ISBN 0-8194-1546-4.

URL ftp://wuarchive.wustl.
edu/doc/techreports/wustl.
edu/math/papers/rovinj.ps.Z.

Wickerhauser:1994:EWA

[Wic94h] Mladen Victor Wickerhauser.

Wavelets: Algorithms and Applications by Yves Meyer. SIAM Review, 36(3):526–528, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). Book review.

wuarchive:1991:WUS

[wua91] wuarchive. Washington University in St. Louis. Inter-Net Anonymous File Transfer

(ftp) Site wuarchive.wustl.edu
[128.252.135.4], 1991.

Wesfreid:1992:ALT

[WW92] Eva Wesfreid and Mladen Victor Wickerhauser. Adapted local trigonometric transform and speech processing. Preprint 9236, CEREMADE–University of Paris–Dauphine, 1992. URL ftp:

//wuarchive.wustl.edu/doc/
techreports/wustl.edu/math/
papers/alttasp.ps.Z. Submit-

ted to IEEE Transactions on Signal Processing.

Wesfreid:1993:ALT

[WW93] Eva Wesfreid and Mladen Victor Wickerhauser. Adapted local trigonometric transform and speech processing. IEEE Transactions on Signal Processing, 41(12):3596–3600, December 1993. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic). URL ftp:

//wuarchive.wustl.edu/doc/
techreports/wustl.edu/math/
papers/alttasp.ps.Z.

Wesfreid:1994:TPP

[WW94] Eva Wesfreid and Mladen Victor Wickerhauser. Traitement de la parole par ondelettes de Malvar. In J. P. Haton, editor, Reconnaisance Automatique de la Parole, Actes du Séminaire. CRIN/INRIA–Nancy, 10–11 March 1994.

Zubair:1991:CTD

[ZSW91] Lareef Zubair, Kannan R. Sreenivasan, and Mladen Victor Wickerhauser. Compression of turbulence data and images using wavelet packets. In T. Gadsky, S. Sirkar, and C. Speziale, editors, Studies in Turbulence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991.